OpenStudy (anonymous):

The sum of two numbers is 100. The second is three times the first. Find the two numbers.

OpenStudy (eyust707):

"The sum of two numbers is 100" x + y = 100

OpenStudy (eyust707):

the second is three times the first y = 3x

OpenStudy (eyust707):

Substitute "y" for "3x" x+ 3x = 100

OpenStudy (eyust707):

simplify 4x = 100 x = 25

OpenStudy (eyust707):

y = 3x y = 3*25 y=75